### v1

# What Do You See? Enhancing Zero-Shot Image Classification with Multimodal Large Language Models

Abstract
Large language models (LLMs) has been effectively used for many computer vision tasks, including image classification. In this paper, we present a simple yet effective approach for zero-shot image classification using multimodal LLMs. By employing multimodal LLMs, we generate comprehensive textual representations from input images. These textual representations are then utilized to generate fixed-dimensional features in a cross-modal embedding space. Subsequently, these features are fused together to perform zero-shot classification using a linear classifier. Our method does not require prompt engineering for each dataset; instead, we use a single, straightforward, set of prompts across all datasets. We evaluated our method on several datasets, and our results demonstrate its remarkable effectiveness, surpassing benchmark accuracy on multiple datasets. On average over ten benchmarks, our method achieved an accuracy gain of 4.1 percentage points, with an increase of 6.8 percentage points on the ImageNet dataset, compared to prior methods. Our findings highlight the potential of multimodal LLMs to enhance computer vision tasks such as zero-shot image classification, offering a significant improvement over traditional methods.
1 Introduction
Zero-shot image classification is a challenging computer vision task where the objective is to classify images into categories that were not seen during method development (i.e., model training for learning-based methods). Recent learning-based approaches for zero-shot image classification (e.g., [1]) utilize large language models (LLMs), such as GPT-4 [2], to generate prompts (i.e., captions or descriptions) for the target classes.
During classification, a vision-language embedding model, such as CLIP [3], is employed to generate features from the input image and the target class prompts, where all input and target features reside in a common embedding space. Subsequently, the input image features are matched to the most similar target class feature, and that class is predicted as the final output. Despite the practicality of this approach, relying solely on visual features from the image during inference may limit classification accuracy. This limitation arises from the potential inability of visual features to capture fine descriptive details present in the class prompts.
Figure 1: An illustration of the difference between our work and prior work (e.g., CuPL [1]) in terms of using LLMs. Prior works use LLMs to describe class labels while we use multimodal LLMs to describe input images and class labels as well as making initial predictions. The shown image is sourced from the ImageNet dataset [4].
In this work, we tackle this limitation by extracting more representative features from the input image. Unlike previous methods, we leverage multimodal LLMs, such as Gemini [5], during the inference phase to generate rich and finely detailed textual representations (e.g., image descriptions) from the input images (refer to Figure 1).
To generate these textual representations, we utilize a straightforward set of prompts that are applied uniformly across all datasets. Consequently, we circumvent the need for extensive prompt engineering or the creation of distinct prompts for each dataset, as required in prior work (e.g., [3, 6]). The textual representations generated by the multimodal LLM are subsequently fed into a contrastive vision-language model (e.g., [3]) to produce features in a cross-modal embedding space. The resulting features are then combined with the visual image features to conduct zero-shot image classification. By employing multimodal LLMs and a straightforward set of prompts, our method outperforms previous zero-shot image classification methods on ten image classification benchmark datasets, achieving an average accuracy gain of 4.1 percentage points and an accuracy increase of 6.8% on the ImageNet dataset [4].
Contribution  The contribution of this paper can be summarized as follows. We introduce a novel method for zero-shot image classification that overcomes the limitations of relying solely on visual image features during inference. Instead, our approach extracts rich textual features directly from input images using multimodal LLMs. These textual features, when combined with visual image features, facilitate zero-shot classification, resulting in significant improvements in classification accuracy. Moreover, our method utilizes a simple set of universal prompts, eliminating the need for dataset-specific prompt engineering. This simplicity makes our method straightforward and easy to implement. As demonstrated through experiments, our approach outperforms prior zero-shot image classification methods across ten benchmark datasets.
The remainder of this paper is organized as follows. In Section 2, we provide a brief overview of prior work related to our method. Section 3 details our proposed approach for zero-shot image classification using LLMs. We then present the experimental results in Section 4, where we validate our method across several benchmark datasets for image classification. Sections 5 and 6 discuss the computational resources and potential limitations of our approach. Finally, we conclude the paper in Section 7.
2 Related work
As our method targets zero-shot image classification, we first provide a brief review of prior work in this area, focusing on methods that utilize open-vocabulary models for the task. For a more comprehensive review of zero-shot image classification, we direct the reader to [7, 8]. Subsequently, as our method relies on multimodal LLMs, we provide a brief summary of previous efforts in the literature regarding the development of such models.
2.1 Zero-shot image classification
Zero-shot image classification has been an active area of research in recent years. Recently, open-vocabulary models (e.g., CLIP [3] and VirTex [9]) have emerged as a promising approach for zero-shot image classification. Such models can accurately classify images across multiple datasets without requiring specific training data for each task. They achieve this by using large amounts of image-text data from the internet to learn how to match images with their corresponding captions. This provides flexibility during image classification, as categories can be defined and modified on the fly without the need for retraining. Instead of traditional classification methods, these models determine image categories by comparing the similarity between an image and a text prompt representing each category. The DeViSE [10] method learns a joint embedding space for images and text, allowing for the transfer of knowledge from seen to unseen classes. CLIP [3] is a contrastive learning-based approach that trains a model to align image and text representations in a shared embedding space. ALIGN [11] proposes a two-stage framework that first aligns image and text embeddings and then refines the classification using a generative model.
2.2 Multimodal large language models
LLMs have been extended to multimodal settings to handle both text and visual data. These models demonstrate impressive capabilities in various tasks. Models like DALL-E [12] and Stable Diffusion [13] can generate descriptive captions for given images. Models like ViLBERT [14] and LXMERT [15] can perform visual question answering, answering questions based on both images and text. Models like GLIDE [16] can generate images from text prompts, providing a flexible way to generate visual content. These multimodal LLMs have found applications in various domains, including content creation, visual search, and accessibility. In this work, we improve the performance of zero-shot image classification by leveraging the outstanding abilities of multimodal LLMs in generating rich and descriptive textual representations from input images.
3 Method
Figure 2: We propose a zero-shot image classification method that leverages multimodal large language models (LLMs) to enhance the accuracy of standard zero-shot classification. Our method employs a set of engineered prompts to generate image description and initial class prediction by the LLM. Subsequently, we encode this data along with the input testing image using a cross-modal embedding encoder to project the inputs into a common feature space. Finally, we fuse the generated features to produce the final query feature, which is then utilized by a standard zero-shot linear image classifier to predict the final class. The shown image is sourced from the Pets dataset [17].
Given an input image, X, containing object(s) belonging to a single class label from a finite set of class labels, $\{l_{i}\}_{i=1}^{m}$, our objective is to classify X without any dataset-specific training process for image classification. The overview of our method is illustrated in Figure 2. As shown in Figure 2, our approach relies on a cross-modal embedding encoder models (image encoder, $f_{i}$, and text encoder, $f_{t}$), trained to learn joint representations of images and text, as demonstrated in prior works, such as [3, 10, 11]. Additionally, we utilize a multimodal LLM, $g$, which is pre-trained on a large corpus of multimodal data. This model, $g$, is designed to generate responses that align with both textual and visual inputs, effectively integrating information from both modalities to make predictions (e.g., [5]).
3.1 Class label features
Our zero-shot classifier includes class label features of the target dataset. We first encode the set of class labels, $\{l_{i}\}_{i=1}^{m}$, in some embedding space by the cross-modal encoder models, $f_{i}$, and $f_{t}$, such that $\{\textbf{L}_{i}\}_{i=1}^{m}$ refers to the set of encoded class label features, where $\textbf{L}_{i}\in\mathbb{R}^{n}$ is a normalized $n$-D class label feature of the class label $l_{i}$ and $n$ is the dimensionality of the embedded features.
Our zero-shot classifier uses the embedded class labels by representing $\{\textbf{L}_{i}\}$ as a 2D matrix, $\textbf{M}\in\mathbb{R}^{n\times m}$, by stacking all encoded class label features:
$\textbf{M}=[\textbf{L}_{1},\dots,\textbf{L}_{m}].$ (1)
Such an encoded feature matrix can be generated using one of three options: (1) directly from the textual class labels, (2) using a human-designed template (e.g., ‘‘A photo of {class_label}[3, 6]), where {class_label} refers to the textual label of one of our classes $\{l_{i}\}_{i=1}^{m}$, or (3) LLM-generated class description(s) [1]. In option (3), the LLM-generated class description(s) are then converted to embedded features, followed by fusion (e.g., averaging) to generate a single embedded feature for each class label in the dataset. Optionally, all features from the three options can be fused together (e.g., averaged) for increased robustness.
Figure 3: Our method utilizes image description and initial class prediction generated by LLM, in addition to the input image, to improve the zero-shot classification accuracy of cross-modal embedding models, such as CLIP [3]. The shown image is from the ImageNet dataset [4].
3.2 Cross-modal input features
To predict the final class, we first encode the input image by the cross-modal image encoder model, $f_{i}$, to generate the image feature $\textbf{X}_{\texttt{if}}\in\mathbb{R}^{n}$ as described below:
$_{\texttt{if}}=f_{i}\left(\textbf{X}\right),$ X ) , (2)
$\textbf{X}_{\texttt{if}}=\frac{1}{\|\widetilde{\textbf{X}}_{\texttt{if}}\|}\widetilde{\textbf{X}}_{\texttt{if}},$ (3)
[…]
$\widetilde{\textbf{X}}_{\texttt{df}}=\left(f_{t}\circ g\right)\left(\textbf{X},p_{\texttt{d}}\right),$ (4)
$\textbf{X}_{\texttt{df}}=\frac{1}{\|\widetilde{\textbf{X}}_{\texttt{df}}\|}\widetilde{\textbf{X}}_{\texttt{df}},$ (5)
$\widetilde{\textbf{X}}_{\texttt{pf}}=\left(f_{t}\circ g\right)\left(\textbf{X},p_{\texttt{c}}\right),$ (6)
$\textbf{X}_{\texttt{pf}}=\frac{1}{\|\widetilde{\textbf{X}}_{\texttt{pf}}\|}\widetilde{\textbf{X}}_{\texttt{pf}},$ (7)
where $\textbf{X}_{\texttt{df}}\in\mathbb{R}^{n}$ and $\textbf{X}_{\texttt{pf}}\in\mathbb{R}^{n}$ refer to the image description feature and initial class prediction feature, respectively. Notably, in our method, unlike prior methods (e.g., [1, 3, 6]), such prompts $p_{\texttt{d}}$ and $p_{\texttt{c}}$ do not require dataset-specific engineering for each dataset. Instead, we employ fixed prompts: the first prompt instructs the LLM to provide a generic image description, while the second prompt includes the textual class labels of the target dataset. Further details regarding our prompts are provided in the supplemental materials (Appendix A).
After generating the three input features (image feature, description feature, and initial prediction feature), we fuse them to generate our final query feature, $\textbf{X}_{\texttt{q}}$. One can interpret this fusion as an ensemble of different candidate features to generate a more precise query feature (see Figure 4). We adopted a simple fusion, where the final query feature, $\textbf{X}_{\texttt{q}}$, is generated by:
$\widetilde{\textbf{X}}_{\texttt{q}}={\textbf{X}_{\texttt{if}}+\textbf{X}_{\texttt{df}}+\textbf{X}_{\texttt{pf}}}\ ,$ (8)
$\textbf{X}_{\texttt{q}}=\frac{1}{\|\widetilde{\textbf{X}}_{\texttt{q}}\|}\widetilde{\textbf{X}}_{\texttt{q}}\ .$ (9)
We found that this simple fusion yields good results compared to alternative fusion approaches. Refer to Section 4.2 for ablation studies.
3.3 Class label prediction
After computing our query feature, $\textbf{X}_{\texttt{q}}$, we apply our zero-shot linear classifier weights, M, to the fused query feature, $\textbf{X}_{\texttt{q}}$, to generate the final similarity scores (i.e., “logits”) of our prediction. This process can be described as follows:
Figure 4: Our query feature is a fusion of features extracted from the input image, image description, and initial prediction. This fusion operates similarly to ensembling, where our fused query feature demonstrates better robustness, achieving higher accuracy compared to traditional image features used in cross-modal-based zero-shot image classification (e.g., CLIP [3]). On the right, the t-SNE plot [18] shows the class-embedded features of the ImageNet dataset [4] (in gray, with the ground-truth class of the shown image in green), our input features (in red), and the query feature after fusion (in blue).
$\textbf{W}=\textbf{X}_{\texttt{q}}^{T}\textbf{M},$ (10)
[…]
4 Experiments
In our experiments, we employed Gemini Pro [5] as our multimodal LLM, $g$, for generating image descriptions and initial predictions. We utilized CLIP (ViT-L/14) [3] as our cross-modal embedding encoder models, $f_{i}$ and $f_{t}$, to encode input testing images, image descriptions, and initial predictions generated by Gemini Pro.
We explored four different versions of class label features, M, in Equation 1. Specifically, we used CLIP to encode the following representations of class labels: 1) class label names, 2) the text template ‘‘A photo of {class_label}’’, where {class_label} denotes each class label in each dataset, and 3) class descriptions, similar to those generated by CuPL [1], and 4) a combination of the aforementioned three options—akin to the fusion of our input features, we combined the three encoded features of each class and computed their average feature. The class descriptions were produced by prompting Gemini Pro to describe each class label in the dataset 50 times, resulting in 50 different class descriptions for each class. Subsequently, we utilized CLIP to encode the 50 class descriptions for each class and compute the average encoded feature vector to represent each class label. The class description features are generated once as described in Section 3. The exact prompts used to generate image descriptions, initial predictions, and class description are detailed in Appendix A.
4.1 Results
We evaluated our method on the following datasets: ImageNet [4], Pets [17], Places365 [19], Food-101 [20], SUN397 [21, 22], Stanford Cars [23], Describable Textures Dataset (DTD) [24], Caltech-101 [25], CIFAR-10 [26], and CIFAR-100 [26]. We compared our results against the following zero-shot classification methods: CLIP [3], SLIP [27], PyramidCLIP [28], nCLIP [29], NLIP [30], UniCLIP [31], ALIP [32], CALIP [6], and CuPL [1]. For CuPL [1], we utilized Gemini Pro [5] for computing class descriptions instead of GPT-3 [2], ensuring a fair comparison with our method, which also employs Gemini Pro [5]. Additionally, it is worth mentioning that the results reported in the original paper of CuPL [1] use different class labels for the ImageNet dataset [4] compared to the original class labels of the dataset. Consequently, we decided to re-compute the results for CuPL using Gemini Pro [5] with the standard ImageNet class labels.
[…]
Dataset (number of classes/number of testing images)
Method
ImageNet
(1K/50K)
C-10
(10/10K)
C-100
(100/10K)
Food
(101/25.3K)
SUN
(397/19.9K)
Cars
(196/8K)
DTD
(47/1.9K)
Caltech
(101/8.7K)
Pets
(37/3.7K)
Places
(365/36.5K)
CLIP (ViT-L/14) [3] 65.1 87.6 54.3 86.8 61.2 65.1 46.7 83.4 87.9 37.3
CLIP (ViT-B/32) [3] 47.0 73.4 41.4 70.8 56.5 43.8 37.0 84.4 72.9 35.6
CLIP (ViT-B/16) [3] 54.9 66.5 37.6 80.1 58.5 52.6 37.9 85.3 80.4 36.3
CLIP (RN50) [3] 43.0 42.7 16.0 57.4 46.3 34.3 30.5 79.1 64.7 29.7
CLIP (RN101) [3] 44.0 49.0 22.4 64.3 50.2 44.5 34.0 83.2 66.3 30.9
SLIP [27] 47.9 87.5 54.2 69.2 56.0 9.0 29.9 80.9 41.6 -
PyramidCLIP [28] 47.8 81.5 53.7 67.8 65.8 65.0 47.2 81.7 83.7 -
nCLIP [29] 48.8 83.4 54.5 65.8 59.9 18.0 57.1 73.9 33.2 -
NLIP [30] 47.4 81.9 47.5 59.2 58.7 7.8 32.9 79.5 39.2 -
UniCLIP [31] 54.2 87.8 56.5 64.6 61.1 19.5 36.6 84.0 69.2 -
ALIP [32] 40.3 83.8 51.9 45.4 47.8 3.4 23.2 74.1 30.7 -
CALIP (ViT-B/32) [6] 60.6 76.5 44.2 77.4 58.6 56.3 42.4 87.7 86.2 36.9
CALIP (ViT-B/16) [6] 57.5 70.2 41.3 80.7 60.6 50.1 39.5 86.2 79.1 38.4
CALIP (RN101) [6] 50.0 49.7 24.5 67.3 51.8 39.4 35.4 84.3 70.5 33.8
CuPL [1] 66.6 86.6 57.7 89.0 65.3 63.9 49.1 90.5 80.0 39.7
Tip-Adapter ($\ast$) [33] 62.0 - - - - - - - - -
SuS-X ($\ast$) [34] 61.9 - - - - - 50.6 - 77.6 -
Tip-Adapter-F ($\diamondsuit$) [33] 65.5 - - - - - - - - -
CLIP-Adapter ($\diamondsuit$) [35] 61.3 - - - - - 66.1 93.4 - -
APE-T ($\diamondsuit$) [36] 66.1 - - - - - - - - -
[…]
Ours (class labels) 69.9 91.8 65.4 92.0 66.8 74.3 53.2 88.5 90.0 41.3
Ours (template) 69.0 93.1 65.8 89.3 64.5 73.6 52.3 84.5 87.9 39.4
Ours (combined) 73.4 93.4 70.2 93.0 70.6 76.6 58.0 89.4 90.9 43.4
For CLIP [3] and CALIP [6], we employed the template “A photo of {class_label}” to encode textual class labels across all datasets, except for Pets [17], DTD [24], and Cars [23] datasets. For these exceptions, we employed specific templates: “A photo of {class_label}, a type of pets”, “A photo of {class_label}, a textural category”, and “A photo of {class_label}, a car model”, respectively. This approach was found to enhance results, consistent with prior findings [3, 37, 38, 39].
For consistency, we resize all images to 224$\times$224 before processing them with our method, CLIP, CuPL, and CALIP methods. Additionally, we report results from training-free and few-shot learning methods for a comprehensive comparison, including: Tip-Adapter [33], SuS-X [34], CLIP-Adapter [35], and APE-T [36].
The top-1 accuracy results are reported in Table 1 (see Appendix B for top-5 accuracy results). As can be seen, our method consistently achieves state-of-the-art results when compared with prior work across all datasets in zero-shot image classification methods, and in the majority of datasets when considering other methods (i.e., training-free and few-shot methods). Using the combined class features yields the most promising results across the majority of datasets, with the exception of Caltech-101 [25], where the best results were achieved using the class description features.
Figure 5: Input data highlighted based on its contribution to the final prediction. Examples are shown from the Places [19] (first row), ImageNet [4] (second row), Food-101 [20] (third row), and CIFAR-100 [26] (last row) datasets.
Figure 5 shows visual examples, where we highlight parts of input modalities (image, initial prediction, and image description) based on their contribution to the final predicted class. To emphasize the significance of each input component in the final prediction, we employ a straightforward approach. Specifically, we utilize a 2D sliding kernel that traverses the image, masking out patches of the image. Subsequently, we measure the difference between the initial prediction and the prediction after masking to highlight areas of the image that contribute most significantly to the final prediction. Similarly, we apply this approach to the textual inputs. Given two distinct input texts – namely, the initial prediction and the image description generated by the LLM (Gemini [5]) – we utilize a sliding kernel with a stride of one word. We mask out words that match the kernel and quantify their importance in our final prediction. As shown in Figure 5, the three inputs collectively contribute to predicting the final class label. In some cases, one or two inputs exhibit a higher level of influence than the others, as demonstrated in the first, second, and third examples.
It is worth mentioning that, while Gemini’s initial predictions do not match the ground-truth class in the first and second examples, the predictions are contextually sensible. In the first example, Gemini’s prediction was ‘tunnel’ which, while not directly matching any class label in the Places dataset [19], conceptually aligns with the displayed ‘catacomb’ image as an underground passage. Similarly, in the second example, Gemini’s initial prediction was ‘Snoopy’, which corresponds to the character drawn on the pen shown in the input image. However, ‘Snoopy’ is not one of the ImageNet [4] class labels and the correct class of the shown image in second row of Figure 5 is ‘ballpoint pen’. This behavior of LLMs is the reason we cannot use them directly as image classifiers, because they sometimes do not restrict the output class to the provided list of target classes. However, such behavior might be beneficial to other classification tasks that are not restricted to a specific set of classes. Additional examples are provided in the supplemental materials (Appendix B).
4.2 Ablation Studies
Table 2: Ablation study on the impact of features used by our method on the classification accuracy. DF refers to the description feature, PF refers to the prediction feature, and IF refers to the image feature. In all datasets, we employed the best class feature as indicated in Table 1. Specifically, we utilized the combined class feature for all datasets except for Caltech-101 [25], where we opted for the class description feature. The best results are highlighted in bold.
Dataset
Method ImageNet C-10 C-100 Food SUN Cars DTD Caltech Pets Places
Ours (DF) 58.6 90.1 64.5 82.7 49.0 65.4 49.8 83.7 48.3 30.1
Ours (PF) 55.7 94.6 73.2 89.6 60.9 70.8 57.7 89.0 87.1 36.1
Ours (DF and PF) 64.5 94.4 74.0 89.8 61.6 71.6 57.7 89.3 87.5 37.0
Ours (DF and IF) 70.7 90.4 64.0 90.8 67.9 71.4 52.6 90.9 85.5 42.1
Ours (PF and IF) 71.6 92.0 67.2 92.2 69.7 74.1 56.4 91.1 90.6 42.7
Ours (DF, PF, and IF) 73.4 93.4 70.2 93.0 70.6 76.6 58.0 91.3 90.9 43.4
We conducted a series of ablation studies to explore different versions of our method and investigate the impact of each feature, different fusion approaches, and different cross-modal embedding models. Table 2 presents the results of our method using solely the encoded image description, referred to as the description feature (DF), as our input. We also report the results obtained by using encoded initial predictions, termed as the prediction feature (PF), as our input, as well as using both DF and PF concurrently as inputs. Additionally, Table 2 shows the results of employing image feature (IF) alongside DF or PF as inputs, and finally, we present the results when leveraging all available inputs – specifically, DF, PF, and IF.
From the results in Table 2, it is clear that incorporating all three features (DF, PF, and IF) yields the best performance across most datasets, except for the CIFAR datasets [26]. This discrepancy may arise from the low resolution of CIFAR images (originally 32$\times$32), where utilizing the IF may degrade accuracy compared to using only DF and PF.
Table 3: Ablation study on various fusion approaches using 5,000 images randomly selected from the ImageNet dataset [4]. The best results are highlighted in bold.
CLIP model [3]
Fusion Approach ViT-L/14 ViT-B/32 ViT-B/16
Max similarity 57.6 57.8 57.9
Avg similarity 66.1 65.5 65.8
Avg feature 72.5 65.7 68.5
Table 3 shows the results of our second set of ablation studies, where we report the results for 5,000 randomly selected images from the ImageNet dataset [4]. We explore the use of different cross-modal embedding models (CLIP [ViT-L/14], CLIP [ViT-B/32], and CLIP [ViT-B/16]) [3], and additionally investigate different fusion approaches. Rather than using the mean feature vector of our input features (DF, PF, and IF), we calculated the similarity between each input feature separately and the dataset class label features. Subsequently, we fused the similarity scores to generate a single similarity score for each class label in each dataset. We explored two fusion methods: averaging and taking the maximum for each class label. As shown, averaging our three features (IF, DF, PF) yields the best results.
5 Computation resources
Our method relies on a multimodal LLM and cross-modal encoders. The cross-modal encoding takes around 15 ms to encode an image or text on an NVIDIA V100 GPU, while the LLM can be accessed through: 1) Cloud API calls, which do not require local resources to load the model, or 2) loading the model locally for processing, which requires an estimated 16 GPUs/TPUs with approximately 256 GB of memory. Each LLM query takes roughly 700 ms to process. The LLM model is the most intensive operation (as discussed in Section 6), but it can be accelerated using multi-threading.
6 Limitations
Our method introduces a new approach by leveraging multimodal LLMs to enhance the accuracy of zero-shot image classification. However, it is important to acknowledge that there are still some limitations inherent in our proposed method. Since our method relies on multiple queries to a multimodal LLM to generate the required features (i.e., DF and PF), there may be potential constraints when running on devices with limited computational power, and it may consume more time compared to other methods. Nevertheless, we believe that advancements in LLMs will lead to models that can run efficiently on lower computational power. This would enable broader accessibility and applicability of such models, such as Gemini [5], GPT [2], and LLaMA [40].
Our method fails in some cases. Figure 6 shows examples of failure cases, where our method misclassify the input image. While the initial prediction and image description features generally enhance classification accuracy, as demonstrated in Table 2, they can sometimes lead to misclassifications. In the first example in Figure 6, both the image description and initial prediction suggest that the image show a ‘lobster’, whereas it actually shows a ‘crayfish’. Similarly, in the second example, the image description lacks specific features of the cat, while the initial prediction suggests the ‘Bengal’ class label, whereas the actual class label is ‘Egyptian Mau’.
Figure 6: Failure examples of our method, where the initial prediction (and the image description in the first example) adversely influenced our final decision. Results are shown for the Caltech-101 dataset [25] (first row) and the Pets dataset [17] (second row).
7 Conclusion
In this work, we introduced a zero-shot image classification method that relies on multimodal large language models (LLMs). Our approach involves using a multimodal LLM to describe the input testing image and make an initial class prediction based on input testing image and the target class label names. Subsequently, we fuse the encoded features of the image description, initial LLM’s prediction, and input testing image to retrieve similar encoded features to class labels from the target dataset. Our method is straightforward and easy to implement, resulting in significant improvements in zero-shot classification accuracy when compared with prior methods in this domain.
Appendix A Additional Details
In the main paper, we presented our method for zero-shot image classification. The inference process of our method is concisely described in Algorithm 1. As part of our method, we employed a set of prompts. Table 4 shows the prompts used for each step discussed in the main paper that employs the LLM (Gemini Pro [5]). Specifically, we detail the prompts used to: 1) conduct zero-shot image classification with Gemini Pro [5], 2) describe a given testing image, and 3) generate class labels descriptions.
The class descriptions were generated using five prompts, as shown in Table 4, with 10 responses generated for each prompt and class, resulting in 50 class descriptions per class label. To encourage diversity in Gemini Pro’s responses, we set the temperature parameter to a high value of 0.99, as done in [1]. An example of implementing our method, including both classifier construction and the inference process, is shown in Code 1.
Table 4: Details of prompts utilized in our work. Each row represents one query task to the LLM. For instance, ‘image classification’ indicates the utilization of LLM to conduct initial zero-shot image classification, which serves as one of the features in our method. The {classes} variable refers to the class labels of the dataset. The {class_label} variable denotes one of the class labels in the given dataset.
Task Prompt
Image classification You are given an image and a list of class labels. Classify the image given the class labels. Answer using a single word if possible. Here are the class labels: {classes}
Image description What do you see? Describe any object precisely, including its type or class.
Class description 1. Describe what a {class_label} looks like in one or two sentences. 2. How can you identify a {class_label} in one or two sentences? 3. What does a {class_label} look like? Respond with one or two sentences. 4. Describe an image from the internet of a {class_label}. Respond with one or two sentences. 5. A short caption of an image of a {class_label}:
Table 5: Additional results on 5,000 images from the ImageNet dataset [4]. Best result is highlighted in yellow
L/14 B/16 B/32 DistilBERT RoBERTa ROUGE-N-F1 ROUGE-F1 Ours
Top-1 52.1 52.0 55.6 43.7 28.8 54.8 54.9 70.2
In Section 4.1, we visualize examples that highlight the important parts of the inputs contributing to the final predicted class label. In the main paper, we described the approach of sequentially masking out patches from the image and comparing the predicted class with the prediction obtained using the entire unmasked image. Similarly, we follow the same approach for text input by sliding a kernel, masking out words, and comparing the predicted class with our original prediction using inputs without any masking. We used a 2D kernel of size 50$\times$50 pixels with a stride of 10 pixels. If there are no highlighted regions in the image due to the small size of the kernel, we enlarge it by 50 until we reach a kernel size of 200$\times$200 pixels.
For the text kernel, we start with a kernel width of 3 words. If none of the words are highlighted, we reduce it by 1 until we use a 1-word kernel sliding over the text. Each prediction was made using the three inputs: the image, initial prediction, and image description, with one of them having masked out patches or words.
Algorithm 1 Performs zero-shot image classification.
Image X, class labels $\{l_{i}\}_{i=1}^{m}$, class label feature matrix M (Equation 1), multimodal LLM $g$, cross-modal encoders $f_{i}$ $f_{t}$, initial class prediction prompt $p_{\texttt{c}}$, image description prompt $p_{\texttt{d}}$ $_{\texttt{if}}=f_{i}\left(\textbf{X}\right)$ X ) $\triangleright$ Image feature $\textbf{X}_{\texttt{if}}=\widetilde{\textbf{X}}_{\texttt{if}}/\|\widetilde{\textbf{X}}_{\texttt{if}}\|$triangleright Vector normalization $\widetilde{\textbf{X}}_{\texttt{df}}=\left(f_{t}\circ g\right)\left(\textbf{X},p_{\texttt{d}}\right)$triangleright Image description feature $\textbf{X}_{\texttt{df}}=\widetilde{\textbf{X}}_{\texttt{df}}/\|\widetilde{\textbf{X}}_{\texttt{df}}\|$triangleright Vector normalization $\widetilde{\textbf{X}}_{\texttt{pf}}=\left(f_{t}\circ g\right)\left(\textbf{X},p_{\texttt{c}}\right)$triangleright Initial class prediction feature $\textbf{X}_{\texttt{pf}}=\widetilde{\textbf{X}}_{\texttt{pf}}/\|\widetilde{\textbf{X}}_{\texttt{pf}}\|$triangleright Vector normalization $\widetilde{\textbf{X}}_{\texttt{q}}={\textbf{X}_{\texttt{if}}+\textbf{X}_{\texttt{df}}+\textbf{X}_{\texttt{pf}}}$triangleright Fused feature $\textbf{X}_{\texttt{q}}=\widetilde{\textbf{X}}_{\texttt{q}}/\|\widetilde{\textbf{X}}_{\texttt{q}}\|$triangleright Vector normalization $\textbf{W}=\textbf{X}_{\texttt{q}}^{T}\textbf{M}$triangleright Similarity scores $x\leftarrow\texttt{argmax}\left(\textbf{W}\right)$triangleright Predicted class index
Predicted class label $l_{x}$ of input image
1import tensorflow as tf
2import cross_modal_encoder as encoder # for example CLIP
3import llm # for example Gemini Pro
4from fixed_prompts import classification_p, description_p, class_ps # see Table 4.
5
6def create_classifer(class_names, k=50):
7 ’’’Constructs zero-shot image classifier.
8 Args:
9 class_names: A list of class names.
10 k: Number of class descriptions to be generated by the LLM.
11 Returns:
12 A zero-shot image classification model.
13 ’’
14 assert k >= len(class_ps)
15 assert k % len(class_ps) == 0
16 weights = []
17 for class_name in class_names:
18 class_name_feature = encoder.encode_text(class_name)
19 template_feature = encoder.encode_text(f"A photo of {class_name}")
20 llm_class_description = tf.zeros((1, encoder.output_feature_length))
21 for _ in range(k // len(class_ps)):
22 for class_p in class_ps:
23 llm_class_feature = llm.process(class_p.format(class_name), temperature=0.99)
24 llm_class_description += encoder.encode_text(llm_class_feature)
25 llm_class_description /= k
26 class_feature = class_name_feature + template_feature + llm_class_description
27 normalized_class_feature = class_feature / tf.norm(class_feature)
28 weights.append(tf.squeeze(normalized_class_feature))
29 model = {"weights": tf.transpose(tf.convert_to_tensor(weights)),
30 "class_names": class_names}
31 return model
32
33
34def classify(image, classifier):
35 ’’’Performs zero-shot image classification.
36 Args:
37 image: Input testing image.
38 classifier: A zero-shot classification model generated by create_classifier function.
39 Returns:
40 Predicted class name.
41 ’’
42 image_feature = encoder.encode_image(image)
43 image_feature /= tf.norm(image_feature)
44 initial_prediction = llm.process([classification_p, image], temperature=0)
45 prediction_feature = encoder.encode_text(initial_prediction)
46 prediction_feature /= tf.norm(prediction_feature)
47 image_description = llm.process([description_p, image], temperature=0)
48 description_feature = encoder.encode_text(image_description)
49 description_feature /= tf.norm(description_feature)
50 query_feature = image_feature + prediction_feature + description_feature
51 query_feature /= tf.norm(query_feature)
52 index = tf.argmax(tf.linalg.matmul(query_feature, classifier["weights"]))
53 return classifier["class_names"][index.numpy().squeeze()]
Code 1: Example Python implementation of our method. In this example, we utilize the combined class feature, as described in Section 4.
Appendix B Additional Results
In this section, we provide supplementary results to those presented in the main paper. Figure 7 shows the confusion matrix for CLIP (ViT-L/14) [3] and our method across two datasets (Caltech-101 [25] and CIFAR-100 [26]). The shown results demonstrate that our method enhances classification accuracy and reduces misclassification rates.
Figure 7: Confusion matrices for zero-shot image classification results of (A) CLIP (ViT-L/14) [3] and (B) our method on the Caltech-101 [25] and CIFAR-100 [26] datasets.
Figure 8: Additional examples demonstrating the influence of input data on final predictions. Examples are provided from the following datasets: SUN397 [21, 22] (first row), DTD [24] (second row), and ImageNet [4] (last three rows).
Table 6: Top-5 classification accuracy of CLIP [3], CuPL [1], and our method on the following datasets: ImageNet [4], CIFAR-10 (C-10) [26], CIFAR-100 (C-100) [26], Food-101 [20], SUN397 [21, 22], Cars [23], DTD [24], Caltech-101 [25], Pets [17], and Places [19]. We report our results with the following class label features: 1) class descriptions, 2) class labels, 3) the template “A photo of {class}”, and 4) combined features of (1-3). The best results are highlighted in yellow.
Dataset
Method ImageNet C-10 C-100 Food SUN Cars DTD Caltech Pets Places
CLIP (ViT-L/14) [3] 88.4 98.5 77.0 97.8 89.1 93.7 72.8 95.2 96.6 64.5
CuPL [1] 91.0 98.1 79.3 98.3 92.1 94.2 77.7 99.8 96.2 68.8
[…]
Ours (class labels) 89.9 99.2 84.3 98.8 91.1 97.5 77.6 98.9 98.6 67.5
Ours (template) 89.8 99.6 84.8 97.9 90.1 97.5 77.2 96.6 97.3 65.3
Ours (combined) 93.0 99.6 88.5 99.0 94.4 97.9 83.8 99.9 99.5 70.9
Table 7: Cohen’s Kappa score of CLIP [3], CuPL [1], and our method on the following datasets: ImageNet [4], CIFAR-10 (C-10) [26], CIFAR-100 (C-100) [26], Food-101 [20], SUN397 [21, 22], Cars [23], DTD [24], Caltech-101 [25], Pets [17], and Places [19]. We report our results with the following class label features: 1) class descriptions, 2) class labels, 3) the template “A photo of {class}”, and 4) combined features of (1-3). The best results are highlighted in yellow.
Dataset
Method ImageNet C-10 C-100 Food SUN Cars DTD Caltech Pets Places
CLIP (ViT-L/14) [3] 0.651 0.862 0.539 0.867 0.611 0.656 0.452 0.830 0.835 0.372
CuPL [1] 0.665 0.851 0.573 0.889 0.652 0.637 0.480 0.902 0.795 0.395
[…]
Ours (class labels) 0.699 0.909 0.650 0.920 0.667 0.742 0.522 0.882 0.897 0.411
Ours (template) 0.709 0.925 0.675 0.919 0.668 0.747 0.545 0.874 0.892 0.412
Ours (combined) 0.734 0.927 0.699 0.929 0.706 0.764 0.571 0.890 0.907 0.432
In Table 2, encouraging results were demonstrated by utilizing the feature of initial prediction produced by the LLM (i.e., Gemini Pro [5]) for zero-shot image classification. Based on these results, one might argue for the direct utilization of Gemini’s class prediction, aiming to match a specific class label from the dataset. However, in several cases, Gemini’s response does not precisely match one of the class labels (as shown in Figure 5). For example, if a ground-truth class label is ‘cat’, Gemini’s response might be ‘The image class is cat’. This discrepancy motivated us to report results of using only Gemini prediction.
In this section, we present additional results from early experiments aimed at utilizing Gemini’s predictions to precisely match one of the class labels in the given dataset. Specifically, we randomly selected 5,000 images from ImageNet [4] for evaluation. While our method, as presented in the main paper, offers a practical way of utilizing Gemini’s predictions, we also present the results of some alternative approaches aimed at precisely identifying one of the dataset class labels, rather than solely relying on the class prediction text generated by Gemini.
Table 5 show the results on the 5,000 images from ImageNet [4] of our main method and alternatives that utilize Gemini’s class prediction to conduct similarity matching with the target dataset class labels. Specifically, we report the results of encoding Gemini’s class prediction using an open-vocabulary language model and measuring the similarity with the encoded class label features. Here, we show the results of using CLIP (ViT-L/14, ViT-B/32, ViT-B/16) [3], DistilBERT [41], and RoBERTa [42].
In addition, we explored classical text similarity metrics – namely, ROUGE-N-F1 and ROUGE-F1 [43] – rather than encoding both Gemini’s prediction and class labels using an open-vocabulary encoding model. As shown in Table 5, our method, which utilizes Gemini’s class prediction as one of the input features, achieves the best results when compared with the alternative approaches.
In the main paper, we reported the top-1 classification accuracy on several datasets [4, 25, 20, 23, 26, 24, 19, 21, 22, 17]. Table 6 presents the top-5 classification accuracy of our method compared to prior work, while Table 7 shows the Cohen’s kappa coefficient. As can be seen, our method achieves a notable improvement while remaining simple and easy to implement.
Lastly, Figure 8 shows additional visual examples, where we highlight the most significant contributors from input parts that influence the final predictions of our method.
Appendix C Broader Impact
Our work introduces a method for zero-shot image classification that leverages the power of multimodal large language models (LLMs) not only during the classifier model construction phase but also at inference time. We achieve this by generating comprehensive textual representations directly from input images. These representations are then combined with the input images for classification, resulting in a significant enhancement in accuracy.
Importantly, our approach eliminates the need for dataset-specific prompt engineering, as commonly required in prior approaches, thereby simplifying the implementation process and enhancing accessibility – effectively acting as a plug-and-play solution. By removing the requirement for dataset-specific customization, our method offers a straightforward and user-friendly approach to zero-shot image classification, making it more accessible to a broader range of users.
By demonstrating its effectiveness across diverse datasets,we illustrate the utility of our method for robust and generalizable real-world computer vision systems reliant on image classification, eliminating the need for dataset-specific training, tuning, or prompt engineering. This approach holds promise for simplifying the deployment of image classification systems and advancing the field of computer vision.

### v2

@@
 Abstract
-Large language models (LLMs) has been effectively used for many computer vision tasks, including image classification.
+Large language models (LLMs) have been effectively used for many computer vision tasks, including image classification.
 In this paper, we present a simple yet effective approach for zero-shot image classification using multimodal LLMs.
@@
 1 Introduction
-Zero-shot image classification is a challenging computer vision task where the objective is to classify images into categories that were not seen during method development (i.e., model training for learning-based methods).
-Recent learning-based approaches for zero-shot image classification (e.g., [1]) utilize large language models (LLMs), such as GPT-4 [2], to generate prompts (i.e., captions or descriptions) for the target classes.
-During classification, a vision-language embedding model, such as CLIP [3], is employed to generate features from the input image and the target class prompts, where all input and target features reside in a common embedding space.
-Subsequently, the input image features are matched to the most similar target class feature, and that class is predicted as the final output.
-Despite the practicality of this approach, relying solely on visual features from the image during inference may limit classification accuracy.
-This limitation arises from the potential inability of visual features to capture fine descriptive details present in the class prompts.
-Figure 1: An illustration of the difference between our work and prior work (e.g., CuPL [1]) in terms of using LLMs.
+Zero-shot image classification aims to categorize images into classes unseen during training, presenting a significant challenge in computer vision.
+Recent approaches leverage the power of large language models (LLMs) like GPT-4 (Brown et al.
+(2020)) to generate prompts for target classes, often in conjunction with vision-language models such as CLIP (Radford et al.
+(2021)) to embed images and text in a common space.
+Open-vocabulary models like CLIP (Radford et al.
+(2021)) and VirTex (Desai & Johnson (2021)) have shown promise in this area due to their ability to generalize to unseen classes.
+These models learn to match images with captions from vast amounts of image-text data, allowing for dynamic classification without retraining.
+Early works like DeViSE (Frome et al.
+(2013)) pioneered the concept of a joint embedding space for images and text, enabling generalization to unseen classes.
+Approaches like CLIP (Radford et al.
+(2021)), based on contrastive learning, and ALIGN (Jia et al.
+(2021)), employing a two-stage framework, further refined the alignment of image and text representations.
+More recent approaches for zero-shot image classification (e.g., Pratt et al.
+(2023); Menon & Vondrick (2023)) have utilized LLMs to generate prompts (i.e., captions or descriptions) for the target classes to further improve the classification accuracy.
+Figure 1: An illustration of the difference between our work and prior work (e.g., CuPL (Pratt et al.
+(2023))) in terms of using LLMs.
 Prior works use LLMs to describe class labels while we use multimodal LLMs to describe input images and class labels as well as making initial predictions.
-The shown image is sourced from the ImageNet dataset [4].
-In this work, we tackle this limitation by extracting more representative features from the input image.
-Unlike previous methods, we leverage multimodal LLMs, such as Gemini [5], during the inference phase to generate rich and finely detailed textual representations (e.g., image descriptions) from the input images (refer to Figure 1).
-To generate these textual representations, we utilize a straightforward set of prompts that are applied uniformly across all datasets.
-Consequently, we circumvent the need for extensive prompt engineering or the creation of distinct prompts for each dataset, as required in prior work (e.g., [3, 6]).
-The textual representations generated by the multimodal LLM are subsequently fed into a contrastive vision-language model (e.g., [3]) to produce features in a cross-modal embedding space.
-The resulting features are then combined with the visual image features to conduct zero-shot image classification.
-By employing multimodal LLMs and a straightforward set of prompts, our method outperforms previous zero-shot image classification methods on ten image classification benchmark datasets, achieving an average accuracy gain of 4.1 percentage points and an accuracy increase of 6.8% on the ImageNet dataset [4].
-Contribution  The contribution of this paper can be summarized as follows.
-We introduce a novel method for zero-shot image classification that overcomes the limitations of relying solely on visual image features during inference.
-Instead, our approach extracts rich textual features directly from input images using multimodal LLMs.
-These textual features, when combined with visual image features, facilitate zero-shot classification, resulting in significant improvements in classification accuracy.
-Moreover, our method utilizes a simple set of universal prompts, eliminating the need for dataset-specific prompt engineering.
-This simplicity makes our method straightforward and easy to implement.
-As demonstrated through experiments, our approach outperforms prior zero-shot image classification methods across ten benchmark datasets.
-The remainder of this paper is organized as follows.
-In Section 2, we provide a brief overview of prior work related to our method.
-Section 3 details our proposed approach for zero-shot image classification using LLMs.
-We then present the experimental results in Section 4, where we validate our method across several benchmark datasets for image classification.
-Sections 5 and 6 discuss the computational resources and potential limitations of our approach.
-Finally, we conclude the paper in Section 7.
-2 Related work
-As our method targets zero-shot image classification, we first provide a brief review of prior work in this area, focusing on methods that utilize open-vocabulary models for the task.
-For a more comprehensive review of zero-shot image classification, we direct the reader to [7, 8].
-Subsequently, as our method relies on multimodal LLMs, we provide a brief summary of previous efforts in the literature regarding the development of such models.
-2.1 Zero-shot image classification
-Zero-shot image classification has been an active area of research in recent years.
-Recently, open-vocabulary models (e.g., CLIP [3] and VirTex [9]) have emerged as a promising approach for zero-shot image classification.
-Such models can accurately classify images across multiple datasets without requiring specific training data for each task.
-They achieve this by using large amounts of image-text data from the internet to learn how to match images with their corresponding captions.
-This provides flexibility during image classification, as categories can be defined and modified on the fly without the need for retraining.
-Instead of traditional classification methods, these models determine image categories by comparing the similarity between an image and a text prompt representing each category.
-The DeViSE [10] method learns a joint embedding space for images and text, allowing for the transfer of knowledge from seen to unseen classes.
-CLIP [3] is a contrastive learning-based approach that trains a model to align image and text representations in a shared embedding space.
-ALIGN [11] proposes a two-stage framework that first aligns image and text embeddings and then refines the classification using a generative model.
-2.2 Multimodal large language models
-LLMs have been extended to multimodal settings to handle both text and visual data.
-These models demonstrate impressive capabilities in various tasks.
-Models like DALL-E [12] and Stable Diffusion [13] can generate descriptive captions for given images.
-Models like ViLBERT [14] and LXMERT [15] can perform visual question answering, answering questions based on both images and text.
-Models like GLIDE [16] can generate images from text prompts, providing a flexible way to generate visual content.
-These multimodal LLMs have found applications in various domains, including content creation, visual search, and accessibility.
-In this work, we improve the performance of zero-shot image classification by leveraging the outstanding abilities of multimodal LLMs in generating rich and descriptive textual representations from input images.
+The shown image is sourced from the ImageNet dataset (Deng et al.
+(2009)).
+However, relying solely on visual features of the input images during inference can limit accuracy, as these features may not be sufficient to fully capture the nuances present in textual descriptions.
+To address this, we propose a novel method that leverages the capabilities of multimodal LLMs to generate rich textual representations of the input images.
+Multimodal LLMs, such as GPT-4 (Brown et al.
+(2020)) and Gemini (Gemini Team Google (2023)), have demonstrated remarkable abilities in various tasks.
+They are capable of processing and integrating information from various sources like text, images, and audio.
+This allows them to perform tasks that were previously impossible, such as generating detailed image descriptions, answering complex visual questions, and even creating realistic images from text.
+Inspired by these advancements, we utilize a straightforward set of prompts to generate detailed textual descriptions of the input images, eliminating the need for complex prompt engineering seen in previous works (e.g., Radford et al.
+(2021); Guo et al.
+(2023)).
+These textual representations are then fused with visual features to perform zero-shot classification.
+See Figure 1.
+Our method offers several key advantages: it significantly improves classification accuracy by incorporating richer textual information extracted directly from the input images; it employs a simple and universal set of prompts, eliminating the need for dataset-specific prompt engineering; and it outperforms existing methods on a variety of benchmark datasets.
+By employing multimodal LLMs and a straightforward set of prompts, our method outperforms previous zero-shot image classification methods.
+Specifically, our method achieves an average accuracy gain of 4.1 percentage points over ten image classification benchmark datasets and an accuracy increase of 6.8% on the ImageNet dataset (Deng et al.
+(2009)).
+In the following sections, we detail our proposed approach for zero-shot image classification using LLMs (Section 2), present experimental results across ten benchmark datasets (Section 3), analyze the computational resources used (Section 4), discuss limitations (Section 5), and conclude with future directions (Section 6).
 2 Method
@@
 Finally, we fuse the generated features to produce the final query feature, which is then utilized by a standard zero-shot linear image classifier to predict the final class.
-The shown image is sourced from the Pets dataset [17].
+The shown image is sourced from the Pets dataset (Parkhi et al.
+(2012)).
 Given an input image, X, containing object(s) belonging to a single class label from a finite set of class labels, $\{l_{i}\}_{i=1}^{m}$, our objective is to classify X without any dataset-specific training process for image classification.
 The overview of our method is illustrated in Figure 2.
-As shown in Figure 2, our approach relies on a cross-modal embedding encoder models (image encoder, $f_{i}$, and text encoder, $f_{t}$), trained to learn joint representations of images and text, as demonstrated in prior works, such as [3, 10, 11].
+As shown in Figure 2, our approach relies on a cross-modal embedding encoder models (image encoder, $f_{i}$, and text encoder, $f_{t}$), trained to learn joint representations of images and text, as demonstrated in prior works, such as Radford et al.
+(2021); Frome et al.
+(2013); Jia et al.
+(2021).
 Additionally, we utilize a multimodal LLM, $g$, which is pre-trained on a large corpus of multimodal data.
-This model, $g$, is designed to generate responses that align with both textual and visual inputs, effectively integrating information from both modalities to make predictions (e.g., [5]).
+This model, $g$, is designed to generate responses that align with both textual and visual inputs, effectively integrating information from both modalities for enhanced predictions (e.g., Gemini Team Google (2023)).
 2.1 Class label features
-Our zero-shot classifier includes class label features of the target dataset.
+Our zero-shot classifier utilizes class label features of the target dataset.
 We first encode the set of class labels, $\{l_{i}\}_{i=1}^{m}$, in some embedding space by the cross-modal encoder models, $f_{i}$, and $f_{t}$, such that $\{\textbf{L}_{i}\}_{i=1}^{m}$ refers to the set of encoded class label features, where $\textbf{L}_{i}\in\mathbb{R}^{n}$ is a normalized $n$-D class label feature of the class label $l_{i}$ and $n$ is the dimensionality of the embedded features.
@@
 $\textbf{M}=[\textbf{L}_{1},\dots,\textbf{L}_{m}].$ (1)
-Such an encoded feature matrix can be generated using one of three options: (1) directly from the textual class labels, (2) using a human-designed template (e.g., ‘‘A photo of {class_label}[3, 6]), where {class_label} refers to the textual label of one of our classes $\{l_{i}\}_{i=1}^{m}$, or (3) LLM-generated class description(s) [1].
+Such an encoded feature matrix can be generated using one of three options: (1) directly from the textual class labels, (2) using a human-designed template – e.g., ‘‘A photo of {class_label}” (Radford et al.
+(2021); Guo et al.
+(2023)), where {class_label} refers to the textual label of one of our classes $\{l_{i}\}_{i=1}^{m}$, or (3) LLM-generated class description(s) (Pratt et al.
+(2023)).
 In option (3), the LLM-generated class description(s) are then converted to embedded features, followed by fusion (e.g., averaging) to generate a single embedded feature for each class label in the dataset.
 Optionally, all features from the three options can be fused together (e.g., averaged) for increased robustness.
-Figure 3: Our method utilizes image description and initial class prediction generated by LLM, in addition to the input image, to improve the zero-shot classification accuracy of cross-modal embedding models, such as CLIP [3].
-The shown image is from the ImageNet dataset [4].
+Figure 3: Our method utilizes image description and initial class prediction generated by LLM, in addition to the input image, to improve the zero-shot classification accuracy of cross-modal embedding models, such as CLIP (Radford et al.
+(2021)).
+The shown image is from the ImageNet dataset (Deng et al.
+(2009)).
 2.2 Cross-modal input features
@@
 where $\textbf{X}_{\texttt{df}}\in\mathbb{R}^{n}$ and $\textbf{X}_{\texttt{pf}}\in\mathbb{R}^{n}$ refer to the image description feature and initial class prediction feature, respectively.
-Notably, in our method, unlike prior methods (e.g., [1, 3, 6]), such prompts $p_{\texttt{d}}$ and $p_{\texttt{c}}$ do not require dataset-specific engineering for each dataset.
+Notably, in our method, unlike prior methods (e.g., Pratt et al.
+(2023); Radford et al.
+(2021); Guo et al.
+(2023)), such prompts $p_{\texttt{d}}$ and $p_{\texttt{c}}$ do not require dataset-specific engineering for each dataset.
 Instead, we employ fixed prompts: the first prompt instructs the LLM to provide a generic image description, while the second prompt includes the textual class labels of the target dataset.
@@
 Figure 4: Our query feature is a fusion of features extracted from the input image, image description, and initial prediction.
-This fusion operates similarly to ensembling, where our fused query feature demonstrates better robustness, achieving higher accuracy compared to traditional image features used in cross-modal-based zero-shot image classification (e.g., CLIP [3]).
-On the right, the t-SNE plot [18] shows the class-embedded features of the ImageNet dataset [4] (in gray, with the ground-truth class of the shown image in green), our input features (in red), and the query feature after fusion (in blue).
+This fusion operates similarly to ensembling, where our fused query feature demonstrates better robustness, achieving higher accuracy compared to traditional image features used in cross-modal-based zero-shot image classification; e.g., CLIP (Radford et al.
+(2021)).
+On the right, the t-SNE (Van der Maaten & Hinton (2008)) plot shows the class-embedded features of the ImageNet dataset (Deng et al.
+(2009)) (in gray, with the ground-truth class of the shown image in green), our input features (in red), and the query feature after fusion (in blue).
 $\textbf{W}=\textbf{X}_{\texttt{q}}^{T}\textbf{M},$ (10)
@@
 3 Experiments
-In our experiments, we employed Gemini Pro [5] as our multimodal LLM, $g$, for generating image descriptions and initial predictions.
-We utilized CLIP (ViT-L/14) [3] as our cross-modal embedding encoder models, $f_{i}$ and $f_{t}$, to encode input testing images, image descriptions, and initial predictions generated by Gemini Pro.
+In our experiments, we employed Gemini Pro (Gemini Team Google (2023)) as our multimodal LLM, $g$, for generating image descriptions and initial predictions.
+We utilized CLIP (ViT-L/14) (Radford et al.
+(2021)) as our cross-modal embedding encoder models, $f_{i}$ and $f_{t}$, to encode input testing images, image descriptions, and initial predictions generated by Gemini Pro.
 We explored four different versions of class label features, M, in Equation 1.
-Specifically, we used CLIP to encode the following representations of class labels: 1) class label names, 2) the text template ‘‘A photo of {class_label}’’, where {class_label} denotes each class label in each dataset, and 3) class descriptions, similar to those generated by CuPL [1], and 4) a combination of the aforementioned three options—akin to the fusion of our input features, we combined the three encoded features of each class and computed their average feature.
+Specifically, we used CLIP to encode the following representations of class labels: 1) class label names, 2) the text template ‘‘A photo of {class_label}’’, where {class_label} denotes each class label in each dataset, and 3) class descriptions, similar to those generated by CuPL (Pratt et al.
+(2023)), and 4) a combination of the aforementioned three options—akin to the fusion of our input features, we combined the three encoded features of each class and computed their average feature.
 The class descriptions were produced by prompting Gemini Pro to describe each class label in the dataset 50 times, resulting in 50 different class descriptions for each class.
@@
 3.1 Results
-We evaluated our method on the following datasets: ImageNet [4], Pets [17], Places365 [19], Food-101 [20], SUN397 [21, 22], Stanford Cars [23], Describable Textures Dataset (DTD) [24], Caltech-101 [25], CIFAR-10 [26], and CIFAR-100 [26].
-We compared our results against the following zero-shot classification methods: CLIP [3], SLIP [27], PyramidCLIP [28], nCLIP [29], NLIP [30], UniCLIP [31], ALIP [32], CALIP [6], and CuPL [1].
-For CuPL [1], we utilized Gemini Pro [5] for computing class descriptions instead of GPT-3 [2], ensuring a fair comparison with our method, which also employs Gemini Pro [5].
-Additionally, it is worth mentioning that the results reported in the original paper of CuPL [1] use different class labels for the ImageNet dataset [4] compared to the original class labels of the dataset.
-Consequently, we decided to re-compute the results for CuPL using Gemini Pro [5] with the standard ImageNet class labels.
+We evaluated our method on the following datasets: ImageNet (Deng et al.
+(2009)), Pets (Parkhi et al.
+(2012)), Places365 (Zhou et al.
+(2017)), Food-101 (Bossard et al.
+(2014)), SUN397 (Xiao et al.
+(2010; 2016)), Stanford Cars (Krause et al.
+(2013)), Describable Textures Dataset (DTD) (Cimpoi et al.
+(2014)), Caltech-101 (Fei-Fei et al.
+(2004)), CIFAR-10 (Krizhevsky et al.
+(2009)), and CIFAR-100 (Krizhevsky et al.
+(2009)).
+We compared our results against the following zero-shot classification methods: CLIP (Radford et al.
+(2021)), SLIP (Mu et al.
+(2022)), PyramidCLIP (Gao et al.
+(2022)), nCLIP (Zhou et al.
+(2023)), NLIP (Huang et al.
+(2023)), UniCLIP (Lee et al.
+(2022)), ALIP (Yang et al.
+(2023)), CALIP (Guo et al.
+(2023)), and CuPL (Pratt et al.
+(2023)).
+For CuPL (Pratt et al.
+(2023)), we utilized Gemini Pro (Gemini Team Google (2023)) for computing class descriptions instead of GPT-3 (Brown et al.
+(2020)), ensuring a fair comparison with our method, which also employs Gemini Pro.
+Additionally, it is worth mentioning that the results reported in the original paper of CuPL (Pratt et al.
+(2023)) use different class labels for the ImageNet dataset (Deng et al.
+(2009)) compared to the original class labels of the dataset.
+Consequently, we decided to re-compute the results for CuPL using Gemini Pro with the standard ImageNet class labels.
 […]
@@
 (365/36.5K)
-CLIP (ViT-L/14) [3] 65.1 87.6 54.3 86.8 61.2 65.1 46.7 83.4 87.9 37.3
-CLIP (ViT-B/32) [3] 47.0 73.4 41.4 70.8 56.5 43.8 37.0 84.4 72.9 35.6
-CLIP (ViT-B/16) [3] 54.9 66.5 37.6 80.1 58.5 52.6 37.9 85.3 80.4 36.3
-CLIP (RN50) [3] 43.0 42.7 16.0 57.4 46.3 34.3 30.5 79.1 64.7 29.7
-CLIP (RN101) [3] 44.0 49.0 22.4 64.3 50.2 44.5 34.0 83.2 66.3 30.9
-SLIP [27] 47.9 87.5 54.2 69.2 56.0 9.0 29.9 80.9 41.6 -
-PyramidCLIP [28] 47.8 81.5 53.7 67.8 65.8 65.0 47.2 81.7 83.7 -
-nCLIP [29] 48.8 83.4 54.5 65.8 59.9 18.0 57.1 73.9 33.2 -
-NLIP [30] 47.4 81.9 47.5 59.2 58.7 7.8 32.9 79.5 39.2 -
-UniCLIP [31] 54.2 87.8 56.5 64.6 61.1 19.5 36.6 84.0 69.2 -
-ALIP [32] 40.3 83.8 51.9 45.4 47.8 3.4 23.2 74.1 30.7 -
-CALIP (ViT-B/32) [6] 60.6 76.5 44.2 77.4 58.6 56.3 42.4 87.7 86.2 36.9
-CALIP (ViT-B/16) [6] 57.5 70.2 41.3 80.7 60.6 50.1 39.5 86.2 79.1 38.4
-CALIP (RN101) [6] 50.0 49.7 24.5 67.3 51.8 39.4 35.4 84.3 70.5 33.8
-CuPL [1] 66.6 86.6 57.7 89.0 65.3 63.9 49.1 90.5 80.0 39.7
-Tip-Adapter ($\ast$) [33] 62.0 - - - - - - - - -
-SuS-X ($\ast$) [34] 61.9 - - - - - 50.6 - 77.6 -
-Tip-Adapter-F ($\diamondsuit$) [33] 65.5 - - - - - - - - -
-CLIP-Adapter ($\diamondsuit$) [35] 61.3 - - - - - 66.1 93.4 - -
-APE-T ($\diamondsuit$) [36] 66.1 - - - - - - - - -
+CLIP (ViT-L/14) (Radford et al. (2021)) 65.1 87.6 54.3 86.8 61.2 65.1 46.7 83.4 87.9 37.3
+CLIP (ViT-B/32) (Radford et al. (2021)) 47.0 73.4 41.4 70.8 56.5 43.8 37.0 84.4 72.9 35.6
+CLIP (ViT-B/16) (Radford et al. (2021)) 54.9 66.5 37.6 80.1 58.5 52.6 37.9 85.3 80.4 36.3
+CLIP (RN50) (Radford et al. (2021)) 43.0 42.7 16.0 57.4 46.3 34.3 30.5 79.1 64.7 29.7
+CLIP (RN101) (Radford et al. (2021)) 44.0 49.0 22.4 64.3 50.2 44.5 34.0 83.2 66.3 30.9
+SLIP (Mu et al. (2022)) 47.9 87.5 54.2 69.2 56.0 9.0 29.9 80.9 41.6 -
+PyramidCLIP (Gao et al. (2022)) 47.8 81.5 53.7 67.8 65.8 65.0 47.2 81.7 83.7 -
+nCLIP (Zhou et al. (2023)) 48.8 83.4 54.5 65.8 59.9 18.0 57.1 73.9 33.2 -
+NLIP (Huang et al. (2023)) 47.4 81.9 47.5 59.2 58.7 7.8 32.9 79.5 39.2 -
+UniCLIP (Lee et al. (2022)) 54.2 87.8 56.5 64.6 61.1 19.5 36.6 84.0 69.2 -
+ALIP (Yang et al. (2023)) 40.3 83.8 51.9 45.4 47.8 3.4 23.2 74.1 30.7 -
+CALIP (ViT-B/32) (Guo et al. (2023)) 60.6 76.5 44.2 77.4 58.6 56.3 42.4 87.7 86.2 36.9
+CALIP (ViT-B/16) (Guo et al. (2023)) 57.5 70.2 41.3 80.7 60.6 50.1 39.5 86.2 79.1 38.4
+CALIP (RN101) (Guo et al. (2023)) 50.0 49.7 24.5 67.3 51.8 39.4 35.4 84.3 70.5 33.8
+CuPL (Pratt et al. (2023)) 66.6 86.6 57.7 89.0 65.3 63.9 49.1 90.5 80.0 39.7
+Tip-Adapter ($\ast$) (Zhang et al. (2022)) 62.0 - - - - - - - - -
+SuS-X ($\ast$) (Udandarao et al. (2023)) 61.9 - - - - - 50.6 - 77.6 -
+Tip-Adapter-F ($\diamondsuit$) (Zhang et al. (2022)) 65.5 - - - - - - - - -
+CLIP-Adapter ($\diamondsuit$) (Gao et al. (2024)) 61.3 - - - - - 66.1 93.4 - -
+APE-T ($\diamondsuit$) (Zhu et al. (2023)) 66.1 - - - - - - - - -
 […]
@@
 Ours (combined) 73.4 93.4 70.2 93.0 70.6 76.6 58.0 89.4 90.9 43.4
-For CLIP [3] and CALIP [6], we employed the template “A photo of {class_label}” to encode textual class labels across all datasets, except for Pets [17], DTD [24], and Cars [23] datasets.
+For CLIP (Radford et al.
+(2021)) and CALIP (Guo et al.
+(2023)), we employed the template “A photo of {class_label}” to encode textual class labels across all datasets, except for Pets (Parkhi et al.
+(2012)), DTD (Cimpoi et al.
+(2014)), and Cars (Krause et al.
+(2013)) datasets.
 For these exceptions, we employed specific templates: “A photo of {class_label}, a type of pets”, “A photo of {class_label}, a textural category”, and “A photo of {class_label}, a car model”, respectively.
-This approach was found to enhance results, consistent with prior findings [3, 37, 38, 39].
+This approach was found to enhance results, consistent with prior findings (Radford et al.
+(2021); Li et al.
+(2023); Allingham et al.
+(2023); Popp et al.
+(2024)).
 For consistency, we resize all images to 224$\times$224 before processing them with our method, CLIP, CuPL, and CALIP methods.
-Additionally, we report results from training-free and few-shot learning methods for a comprehensive comparison, including: Tip-Adapter [33], SuS-X [34], CLIP-Adapter [35], and APE-T [36].
+Additionally, we report results from training-free and few-shot learning methods for a comprehensive comparison, including: Tip-Adapter (Zhang et al.
+(2022)), SuS-X (Udandarao et al.
+(2023)), CLIP-Adapter (Gao et al.
+(2024)), and APE-T (Zhu et al.
+(2023)).
 The top-1 accuracy results are reported in Table 1 (see Appendix B for top-5 accuracy results).
 As can be seen, our method consistently achieves state-of-the-art results when compared with prior work across all datasets in zero-shot image classification methods, and in the majority of datasets when considering other methods (i.e., training-free and few-shot methods).
-Using the combined class features yields the most promising results across the majority of datasets, with the exception of Caltech-101 [25], where the best results were achieved using the class description features.
+Using the combined class features yields the most promising results across the majority of datasets, with the exception of Caltech-101 (Fei-Fei et al.
+(2004)), where the best results were achieved using the class description features.
 Figure 5: Input data highlighted based on its contribution to the final prediction.
-Examples are shown from the Places [19] (first row), ImageNet [4] (second row), Food-101 [20] (third row), and CIFAR-100 [26] (last row) datasets.
+Examples are shown from the Places (Zhou et al.
+(2017)) (first row), ImageNet (Deng et al.
+(2009)) (second row), Food-101 (Bossard et al.
+(2014)) (third row), and CIFAR-100 (Krizhevsky et al.
+(2009)) (last row) datasets.
 Figure 5 shows visual examples, where we highlight parts of input modalities (image, initial prediction, and image description) based on their contribution to the final predicted class.
@@
 Similarly, we apply this approach to the textual inputs.
-Given two distinct input texts – namely, the initial prediction and the image description generated by the LLM (Gemini [5]) – we utilize a sliding kernel with a stride of one word.
+Given two distinct input texts – namely, the initial prediction and the image description generated by the LLM (Gemini Team Google (2023)) – we utilize a sliding kernel with a stride of one word.
 We mask out words that match the kernel and quantify their importance in our final prediction.
@@
 It is worth mentioning that, while Gemini’s initial predictions do not match the ground-truth class in the first and second examples, the predictions are contextually sensible.
-In the first example, Gemini’s prediction was ‘tunnel’ which, while not directly matching any class label in the Places dataset [19], conceptually aligns with the displayed ‘catacomb’ image as an underground passage.
+In the first example, Gemini’s prediction was ‘tunnel’ which, while not directly matching any class label in the Places dataset (Zhou et al.
+(2017)), conceptually aligns with the displayed ‘catacomb’ image as an underground passage.
 Similarly, in the second example, Gemini’s initial prediction was ‘Snoopy’, which corresponds to the character drawn on the pen shown in the input image.
-However, ‘Snoopy’ is not one of the ImageNet [4] class labels and the correct class of the shown image in second row of Figure 5 is ‘ballpoint pen’.
+However, ‘Snoopy’ is not one of the ImageNet (Deng et al.
+(2009)) class labels and the correct class of the shown image in second row of Figure 5 is ‘ballpoint pen’.
 This behavior of LLMs is the reason we cannot use them directly as image classifiers, because they sometimes do not restrict the output class to the provided list of target classes.
@@
 In all datasets, we employed the best class feature as indicated in Table 1.
-Specifically, we utilized the combined class feature for all datasets except for Caltech-101 [25], where we opted for the class description feature.
+Specifically, we utilized the combined class feature for all datasets except for Caltech-101 (Fei-Fei et al.
+(2004)), where we opted for the class description feature.
 The best results are highlighted in bold.
@@
 Additionally, Table 2 shows the results of employing image feature (IF) alongside DF or PF as inputs, and finally, we present the results when leveraging all available inputs – specifically, DF, PF, and IF.
-From the results in Table 2, it is clear that incorporating all three features (DF, PF, and IF) yields the best performance across most datasets, except for the CIFAR datasets [26].
+From the results in Table 2, it is clear that incorporating all three features (DF, PF, and IF) yields the best performance across most datasets, except for the CIFAR datasets (Krizhevsky et al.
+(2009)).
 This discrepancy may arise from the low resolution of CIFAR images (originally 32$\times$32), where utilizing the IF may degrade accuracy compared to using only DF and PF.
-Table 3: Ablation study on various fusion approaches using 5,000 images randomly selected from the ImageNet dataset [4]. The best results are highlighted in bold.
-CLIP model [3]
+Table 3: Ablation study on various fusion approaches using 5,000 images randomly selected from the ImageNet dataset (Deng et al. (2009)). The best results are highlighted in bold.
+CLIP model (Radford et al. (2021))
 Fusion Approach ViT-L/14 ViT-B/32 ViT-B/16
@@
 Avg feature 72.5 65.7 68.5
-Table 3 shows the results of our second set of ablation studies, where we report the results for 5,000 randomly selected images from the ImageNet dataset [4].
-We explore the use of different cross-modal embedding models (CLIP [ViT-L/14], CLIP [ViT-B/32], and CLIP [ViT-B/16]) [3], and additionally investigate different fusion approaches.
+Table 3 shows the results of our second set of ablation studies, where we report the results for 5,000 randomly selected images from the ImageNet dataset (Deng et al.
+(2009)).
+We explore the use of different cross-modal embedding models (CLIP [ViT-L/14], CLIP [ViT-B/32], and CLIP [ViT-B/16]) (Radford et al.
+(2021)), and additionally investigate different fusion approaches.
 Rather than using the mean feature vector of our input features (DF, PF, and IF), we calculated the similarity between each input feature separately and the dataset class label features.
@@
 Nevertheless, we believe that advancements in LLMs will lead to models that can run efficiently on lower computational power.
-This would enable broader accessibility and applicability of such models, such as Gemini [5], GPT [2], and LLaMA [40].
+This would enable broader accessibility and applicability of such models, such as Gemini (Gemini Team Google (2023)), GPT (Brown et al.
+(2020)), and LLaMA (Touvron et al.
+(2023)).
 Our method fails in some cases.
@@
 Figure 6: Failure examples of our method, where the initial prediction (and the image description in the first example) adversely influenced our final decision.
-Results are shown for the Caltech-101 dataset [25] (first row) and the Pets dataset [17] (second row).
+Results are shown for the Caltech-101 dataset (Fei-Fei et al.
+(2004)) (first row) and the Pets dataset (Parkhi et al.
+(2012)) (second row).
 6 Conclusion
@@
 As part of our method, we employed a set of prompts.
-Table 4 shows the prompts used for each step discussed in the main paper that employs the LLM (Gemini Pro [5]).
-Specifically, we detail the prompts used to: 1) conduct zero-shot image classification with Gemini Pro [5], 2) describe a given testing image, and 3) generate class labels descriptions.
+Table 4 shows the prompts used for each step discussed in the main paper that employs the LLM (Gemini Pro (Gemini Team Google (2023))).
+Specifically, we detail the prompts used to: 1) conduct zero-shot image classification with Gemini Pro (Gemini Team Google (2023)), 2) describe a given testing image, and 3) generate class labels descriptions.
 The class descriptions were generated using five prompts, as shown in Table 4, with 10 responses generated for each prompt and class, resulting in 50 class descriptions per class label.
-To encourage diversity in Gemini Pro’s responses, we set the temperature parameter to a high value of 0.99, as done in [1].
+To encourage diversity in Gemini Pro’s responses, we set the temperature parameter to a high value of 0.99, as done in Pratt et al.
+(2023).
 An example of implementing our method, including both classifier construction and the inference process, is shown in Code 1.
@@
 The {classes} variable refers to the class labels of the dataset.
+The {predicted_class} refers to Gemini Pro’s output of the image classification prompt.
 The {class_label} variable denotes one of the class labels in the given dataset.
@@
 A short caption of an image of a {class_label}:
-Table 5: Additional results on 5,000 images from the ImageNet dataset [4]. Best result is highlighted in yellow
+Table 5: Additional results on 5,000 images from the ImageNet dataset (Deng et al. (2009)). Best result is highlighted in yellow
 L/14 B/16 B/32 DistilBERT RoBERTa ROUGE-N-F1 ROUGE-F1 Ours
@@
 In this section, we provide supplementary results to those presented in the main paper.
-Figure 7 shows the confusion matrix for CLIP (ViT-L/14) [3] and our method across two datasets (Caltech-101 [25] and CIFAR-100 [26]).
+Figure 7 shows the confusion matrix for CLIP (ViT-L/14) (Radford et al.
+(2021)) and our method across two datasets (Caltech-101 (Fei-Fei et al.
+(2004)) and CIFAR-100 (Krizhevsky et al.
+(2009))).
 The shown results demonstrate that our method enhances classification accuracy and reduces misclassification rates.
-Figure 7: Confusion matrices for zero-shot image classification results of (A) CLIP (ViT-L/14) [3] and (B) our method on the Caltech-101 [25] and CIFAR-100 [26] datasets.
+Figure 7: Confusion matrices for zero-shot image classification results of (A) CLIP (ViT-L/14) (Radford et al.
+(2021)) and (B) our method on the Caltech-101 (Fei-Fei et al.
+(2004)) and CIFAR-100 (Krizhevsky et al.
+(2009)) datasets.
 Figure 8: Additional examples demonstrating the influence of input data on final predictions.
-Examples are provided from the following datasets: SUN397 [21, 22] (first row), DTD [24] (second row), and ImageNet [4] (last three rows).
-Table 6: Top-5 classification accuracy of CLIP [3], CuPL [1], and our method on the following datasets: ImageNet [4], CIFAR-10 (C-10) [26], CIFAR-100 (C-100) [26], Food-101 [20], SUN397 [21, 22], Cars [23], DTD [24], Caltech-101 [25], Pets [17], and Places [19].
+Examples are provided from the following datasets: SUN397 (Xiao et al.
+(2010; 2016)) (first row), DTD (Cimpoi et al.
+(2014)) (second row), and ImageNet (Deng et al.
+(2009)) (last three rows).
+Table 6: Top-5 classification accuracy of CLIP (Radford et al.
+(2021)), CuPL (Pratt et al.
+(2023)), and our method on the following datasets: ImageNet (Deng et al.
+(2009)), CIFAR-10 (C-10) (Krizhevsky et al.
+(2009)), CIFAR-100 (C-100) (Krizhevsky et al.
+(2009)), Food-101 (Bossard et al.
+(2014)), SUN397 (Xiao et al.
+(2010; 2016)), Cars (Krause et al.
+(2013)), DTD (Cimpoi et al.
+(2014)), Caltech-101 (Fei-Fei et al.
+(2004)), Pets (Parkhi et al.
+(2012)), and Places (Zhou et al.
+(2017)).
 We report our results with the following class label features: 1) class descriptions, 2) class labels, 3) the template “A photo of {class}”, and 4) combined features of (1-3).
@@
 Method ImageNet C-10 C-100 Food SUN Cars DTD Caltech Pets Places
-CLIP (ViT-L/14) [3] 88.4 98.5 77.0 97.8 89.1 93.7 72.8 95.2 96.6 64.5
-CuPL [1] 91.0 98.1 79.3 98.3 92.1 94.2 77.7 99.8 96.2 68.8
+CLIP (ViT-L/14) (Radford et al. (2021)) 88.4 98.5 77.0 97.8 89.1 93.7 72.8 95.2 96.6 64.5
+CuPL (Pratt et al. (2023)) 91.0 98.1 79.3 98.3 92.1 94.2 77.7 99.8 96.2 68.8
 […]
@@
 Ours (combined) 93.0 99.6 88.5 99.0 94.4 97.9 83.8 99.9 99.5 70.9
-Table 7: Cohen’s Kappa score of CLIP [3], CuPL [1], and our method on the following datasets: ImageNet [4], CIFAR-10 (C-10) [26], CIFAR-100 (C-100) [26], Food-101 [20], SUN397 [21, 22], Cars [23], DTD [24], Caltech-101 [25], Pets [17], and Places [19].
+Table 7: Cohen’s Kappa score of CLIP (Radford et al.
+(2021)), CuPL (Pratt et al.
+(2023)), and our method on the following datasets: ImageNet (Deng et al.
+(2009)), CIFAR-10 (C-10) (Krizhevsky et al.
+(2009)), CIFAR-100 (C-100) (Krizhevsky et al.
+(2009)), Food-101 (Bossard et al.
+(2014)), SUN397 (Xiao et al.
+(2010; 2016)), Cars (Krause et al.
+(2013)), DTD (Cimpoi et al.
+(2014)), Caltech-101 (Fei-Fei et al.
+(2004)), Pets (Parkhi et al.
+(2012)), and Places (Zhou et al.
+(2017)).
 We report our results with the following class label features: 1) class descriptions, 2) class labels, 3) the template “A photo of {class}”, and 4) combined features of (1-3).
@@
 Method ImageNet C-10 C-100 Food SUN Cars DTD Caltech Pets Places
-CLIP (ViT-L/14) [3] 0.651 0.862 0.539 0.867 0.611 0.656 0.452 0.830 0.835 0.372
-CuPL [1] 0.665 0.851 0.573 0.889 0.652 0.637 0.480 0.902 0.795 0.395
+CLIP (ViT-L/14) (Radford et al. (2021)) 0.651 0.862 0.539 0.867 0.611 0.656 0.452 0.830 0.835 0.372
+CuPL (Pratt et al. (2023)) 0.665 0.851 0.573 0.889 0.652 0.637 0.480 0.902 0.795 0.395
 […]
@@
 Ours (combined) 0.734 0.927 0.699 0.929 0.706 0.764 0.571 0.890 0.907 0.432
-In Table 2, encouraging results were demonstrated by utilizing the feature of initial prediction produced by the LLM (i.e., Gemini Pro [5]) for zero-shot image classification.
+In Table 2, encouraging results were demonstrated by utilizing the feature of initial prediction produced by the LLM (i.e., Gemini Pro (Gemini Team Google (2023))) for zero-shot image classification.
 Based on these results, one might argue for the direct utilization of Gemini’s class prediction, aiming to match a specific class label from the dataset.
@@
 In this section, we present additional results from early experiments aimed at utilizing Gemini’s predictions to precisely match one of the class labels in the given dataset.
-Specifically, we randomly selected 5,000 images from ImageNet [4] for evaluation.
+Specifically, we randomly selected 5,000 images from ImageNet (Deng et al.
+(2009)) for evaluation.
 While our method, as presented in the main paper, offers a practical way of utilizing Gemini’s predictions, we also present the results of some alternative approaches aimed at precisely identifying one of the dataset class labels, rather than solely relying on the class prediction text generated by Gemini.
-Table 5 show the results on the 5,000 images from ImageNet [4] of our main method and alternatives that utilize Gemini’s class prediction to conduct similarity matching with the target dataset class labels.
+Table 5 show the results on the 5,000 images from ImageNet (Deng et al.
+(2009)) of our main method and alternatives that utilize Gemini’s class prediction to conduct similarity matching with the target dataset class labels.
 Specifically, we report the results of encoding Gemini’s class prediction using an open-vocabulary language model and measuring the similarity with the encoded class label features.
-Here, we show the results of using CLIP (ViT-L/14, ViT-B/32, ViT-B/16) [3], DistilBERT [41], and RoBERTa [42].
-In addition, we explored classical text similarity metrics – namely, ROUGE-N-F1 and ROUGE-F1 [43] – rather than encoding both Gemini’s prediction and class labels using an open-vocabulary encoding model.
+Here, we show the results of using CLIP (ViT-L/14, ViT-B/32, ViT-B/16) (Radford et al.
+(2021)), DistilBERT (Sanh et al.
+(2019)), and RoBERTa (Liu et al.
+(2019)).
+In addition, we explored classical text similarity metrics – namely, ROUGE-N-F1 and ROUGE-F1 (Lin (2004)) – rather than encoding both Gemini’s prediction and class labels using an open-vocabulary encoding model.
 As shown in Table 5, our method, which utilizes Gemini’s class prediction as one of the input features, achieves the best results when compared with the alternative approaches.
-In the main paper, we reported the top-1 classification accuracy on several datasets [4, 25, 20, 23, 26, 24, 19, 21, 22, 17].
+In the main paper, we reported the top-1 classification accuracy on several datasets (Deng et al.
+(2009); Fei-Fei et al.
+(2004); Bossard et al.
+(2014); Krause et al.
+(2013); Krizhevsky et al.
+(2009); Cimpoi et al.
+(2014); Zhou et al.
+(2017); Xiao et al.
+(2010; 2016); Parkhi et al.
+(2012)).
 Table 6 presents the top-5 classification accuracy of our method compared to prior work, while Table 7 shows the Cohen’s kappa coefficient.